# Jacobian Descent for Multi-Objective Optimization

### v2

@@ 1 Introduction
The field of multi-objective optimization studies minimization of vector-valued objective functions (Sawaragi et al., 1985; Ehrgott, 2005; Branke, 2008; Deb et al., 2016). In deep learning, a widespread approach to train a model with multiple objectives is to combine those into a scalar loss function minimized by stochastic gradient descent. While this method is simple, it comes at the expense of potentially degrading some individual objectives. Without prior knowledge of their relative importance, this is undesirable.
Early works have attempted to extend gradient descent (GD) to consider several objectives simultaneously, and thus several gradients (Fliege & Svaiter, 2000; Désidéri, 2012). Essentially, they propose a heuristic to prevent the degradation of any individual objective. Several other works have built upon this method, analyzing its convergence properties or extending it to a stochastic setting (Fliege et al., 2019; Poirion et al., 2017; Mercier et al., 2018). Later, this has been applied to multi-task learning to tackle conflict between tasks, illustrated by contradicting gradient directions (Sener & Koltun, 2018). Many studies have followed, proposing various other algorithms for the training of multi-task models (Yu et al., 2020; Liu et al., 2021a, b; Lin et al., 2021; Navon et al., 2022; Senushkin et al., 2023; Chen et al., 2020). They commonly rely on an aggregator that maps a collection of task-specific gradients (a Jacobian matrix) to a shared parameter update.
@@ 2.1 Jacobian descent
where $o(\|{\bm{y}}\|)$ indicates that $\lim_{\|{\bm{y}}\|\to 0}\frac{{\bm{f}}({\bm{x}}+{\bm{y}})-{\bm{f}}({\bm{x}})-\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}}{\|{\bm{y}}\|}={\bm{0}}$. The term ${\bm{f}}({\bm{x}})+\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$ is the first-order Taylor approximation of ${\bm{f}}({\bm{x}}+{\bm{y}})$. Via this approximation, we aim to select a small update ${\bm{y}}$ that reduces ${\bm{f}}({\bm{x}}+{\bm{y}})$, ideally achieving ${\bm{f}}({\bm{x}}+{\bm{y}})\preceq{\bm{f}}({\bm{x}})$. As the approximation depends on ${\bm{y}}$ only through $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$, selecting the update based on the Jacobian is natural. A mapping $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ reducing such a matrix into a vector is called an aggregator. For any $J\in\mathbb{R}^{m\times n}$, $\mathcal{A}(J)$ is called the aggregation of $J$ by $\mathcal{A}$.
@@ 2.4 Convergence
Another commonly studied type of convergence for multi-objective optimization is convergence to a stationary point. If for a given ${\bm{x}}\in{\mathbb{R}}^{n}$, there exists ${\bm{0}}\precneq{\bm{w}}$ satisfying $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})^{\top}{\bm{w}}=0$ then ${\bm{x}}$ is said to be Pareto stationary. Even though every Pareto optimal point is Pareto stationary, the converse does not hold, even in the convex case. The function $\begin{bmatrix}x&y\end{bmatrix}^{\top}\mapsto\begin{bmatrix}x^{2}&y^{2}\end{bmatrix}^{\top}$ illustrates this discrepancy. Its Pareto set only contains the origin, but its set of Pareto stationary points is the union of the two axes. Despite being necessary, convergence to a Pareto stationary point is thus not a sufficient condition for optimality and, hence, constitutes a rather weak guarantee. To the best of our knowledge, $\mathcal{A}_{\text{UPGrad}}$ is the first non-conflicting aggregator that provably converges to the Pareto front in the smooth convex case.
@@ 4 Existing aggregators
In the context of multi-task learning, several works have proposed iterative optimization algorithms based on the combination of task-specific gradients (Sener & Koltun, 2018; Yu et al., 2020; Liu et al., 2021b, a; Lin et al., 2021; Navon et al., 2022; Senushkin et al., 2023). These methods can be formulated as variants of JD parameterized by different aggregators. More specifically, since the gradients are stochastically estimated from batches of data, these are cases of what we call SEJD. In the following, we briefly present the most prominent aggregators and summarize their properties in Table 1. As a baseline, we also consider $\mathcal{A}_{\text{Mean}}$, which simply averages the rows of the Jacobian. Their formal definitions are provided in Appendix B. Some of them are also illustrated in Figure 1(b).
$\mathcal{A}_{\text{RGW}}$ aggregates the matrix using a random vector of weights (Lin et al., 2021). $\mathcal{A}_{\text{MGDA}}$ gives the aggregation that maximizes the smallest improvement (Désidéri, 2012; Sener & Koltun, 2018; Fliege & Svaiter, 2000). $\mathcal{A}_{\text{CAGrad}}$ maximizes the smallest improvement in a ball around the average gradient whose radius is parameterized by $c\in\left[0,1\right[{}$ (Liu et al., 2021a). $\mathcal{A}_{\text{PCGrad}}$ projects each gradient onto the orthogonal hyperplane of other gradients in case of conflict, iteratively and in a random order (Yu et al., 2020). It is, however, only non-conflicting when $m\varleq 2$, in which case $\mathcal{A}_{\text{PCGrad}}=m\cdot\mathcal{A}_{\text{UPGrad}}$. IMTL-G is a method to balance some gradients with impartiality (Liu et al., 2021b). It is only defined for linearly independent gradients, but we generalize it as a formal aggregator, denoted $\mathcal{A}_{\text{IMTL-G}}$, in Appendix B.6. Aligned-MTL orthonormalizes the Jacobian and weights its rows according to some preferences (Senushkin et al., 2023). We denote by $\mathcal{A}_{\text{Aligned-MTL}}$ this method with uniform preferences. $\mathcal{A}_{\text{Nash-MTL}}$ aggregates Jacobians by finding the Nash equilibrium between task-specific gradients (Navon et al., 2022). Lastly, the GradDrop layer (Chen et al., 2020) defines a custom backward pass that combines gradients with respect to some internal activation. The corresponding aggregator, denoted $\mathcal{A}_{\text{GradDrop}}$, randomly drops out some gradient coordinates based on their sign and sums the remaining ones.
@@
Several other works consider the gradients to be noisy when making their theoretical analysis (Liu & Vicente, 2021; Zhou et al., 2022; Fernando et al., 2022; Chen et al., 2024; Xiao et al., 2024). Their solutions for combining gradients are typically stateful. Unifying their formulations would thus require a more complex variant of Jacobian descent.
In the federated learning setting, several aggregators have been proposed to combine the model updates while being robust to adversaries (Blanchard et al., 2017; Guerraoui et al., 2018; Chen et al., 2017; Yin et al., 2018). We do not study them here as they mainly focus on security aspects.
Table 1: Properties satisfied for any number of objectives. Proofs are provided in Appendix B.
@@ 5 Experiments
We train convolutional neural networks on subsets of SVHN (Netzer et al., 2011), CIFAR-10 (Krizhevsky et al., 2009), EuroSAT (Helber et al., 2019), MNIST (LeCun et al., 1998), Fashion-MNIST (Xiao et al., 2017) and Kuzushiji-MNIST (Clanuwat et al., 2018). To make the comparisons as fair as possible, we have tuned the learning rate very precisely for each aggregator, as explained in detail in Appendix C.1. We have also run the same experiments several times independently to gain confidence in our results. Since this leads to a total of $43776$ training runs across all of our experiments, we have limited the size of each training dataset to 1024 images, greatly reducing computational costs. Note that this is strictly an optimization problem: we are not studying the generalization of the model, which would be captured by some performance metric on a test set. Other experimental settings, such as the network architectures and the total computational budget used to run our experiments, are given in Appendix C. Figure 2 reports the main results on SVHN and CIFAR-10, two of the datasets exhibiting the most substantial performance gap. Results on the other datasets and aggregators are reported in Appendix D.1. They also demonstrate a significant performance gap.
@@
The sub-optimal performance of $\mathcal{A}_{\text{MGDA}}$ in this setting can be attributed to its sensitivity to small gradients. If any row of the Jacobian approaches zero, the aggregation by $\mathcal{A}_{\text{MGDA}}$ will also approach zero. This observation illustrates the discrepancy between stationarity and optimality, as discussed in Section 2.4. A notable advantage of linearity under scaling is to explicitly prevent this from happening.
@@
While increasing the batch size in SGD reduces variance, the effect of doing so in SSJD combined with $\mathcal{A}_{\text{UPGrad}}$ is non-trivial, as it also tightens the dual cone. Additional results obtained when varying the batch size or updating the parameters with the Adam optimizer are available in Appendices D.2 and D.3, respectively.
While an iteration of SSJD is more expensive than an iteration of SGD, its runtime is influenced by several factors, including the choice of aggregator, the parallelization capabilities of the hardware used for Jacobian computation, and the implementation. Appendix E provides an empirical comparison of the runtime of SGD and SSJD at the current state of our library, for various aggregators. Additionally, we propose a path towards a more efficient implementation in the next section.
@@ 7 Conclusion
Our experimentation has some limitations. First, we only evaluate JD on IWRM, a setting with moderately conflicting objectives. It would be essential to develop proper benchmarks to compare aggregators on a wide variety of problems. Ideally, such problems should involve substantially conflicting objectives, e.g. multi-task learning with inherently competing or even adversarial tasks. Then, we have limited our scope to the comparison of optimization speeds, disregarding generalization. While this simplifies the experiments and makes the comparison rigorous, optimization and generalization are sometimes intertwined. We thus believe that future works should focus on both aspects.
@@ Acknowledgments
Appendix A Proofs
A.1 Supplementary theoretical results

### v3

@@ 1 Introduction
The field of multi-objective optimization studies minimization of vector-valued objective functions (Sawaragi et al., 1985; Ehrgott, 2005; Branke, 2008; Deb et al., 2016). In deep learning, a widespread approach to train a model with multiple objectives is to combine those into a scalar loss function minimized by stochastic gradient descent. While this method is simple, it comes at the expense of potentially degrading some individual objectives. Without prior knowledge of their relative importance, this is undesirable. In opposition, multi-objective optimization methods typically attempt to optimize all objectives simultaneously, without making arbitrary compromises: the goal is to find points for which no improvement can be made on some objectives without degrading others.
Early works have attempted to extend gradient descent (GD) to consider several objectives simultaneously, and thus several gradients (Fliege & Svaiter, 2000; Désidéri, 2012). Essentially, they propose a heuristic to prevent the degradation of any individual objective. Several other works have built upon this method, analyzing its convergence properties or extending it to a stochastic setting (Poirion et al., 2017; Mercier et al., 2018; Fliege et al., 2019). Later, this has been applied to multi-task learning to tackle conflict between tasks, illustrated by contradicting gradient directions (Sener & Koltun, 2018). Many studies have followed, proposing various other algorithms for the training of multi-task models (Yu et al., 2020; Chen et al., 2020; Liu et al., 2021a, b; Lin et al., 2021; Navon et al., 2022; Senushkin et al., 2023). They commonly rely on an aggregator that maps a collection of task-specific gradients (a Jacobian matrix) to a shared parameter update.
@@ 2.1 Jacobian descent
The term ${\bm{f}}({\bm{x}})+\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$ is the first-order Taylor approximation of ${\bm{f}}({\bm{x}}+{\bm{y}})$. Via this approximation, we aim to select a small update ${\bm{y}}$ that reduces ${\bm{f}}({\bm{x}}+{\bm{y}})$, ideally achieving ${\bm{f}}({\bm{x}}+{\bm{y}})\preceq{\bm{f}}({\bm{x}})$. As the approximation depends on ${\bm{y}}$ only through $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$, selecting the update based on the Jacobian is natural. A mapping $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ reducing such a matrix into a vector is called an aggregator. For any $J\in\mathbb{R}^{m\times n}$, $\mathcal{A}(J)$ is called the aggregation of $J$ by $\mathcal{A}$.
@@ 2.4 Convergence to the Pareto front
Another common theoretical result for multi-objective optimization is convergence to a weakly stationary point. If for a given ${\bm{x}}\in{\mathbb{R}}^{n}$, there exists ${\bm{0}}\precneq{\bm{w}}$ satisfying $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})^{\top}{\bm{w}}=0$ then ${\bm{x}}$ is said to be weakly Pareto stationary. Even though every Pareto optimal point is weakly Pareto stationary, the converse does not hold, even in the convex case. Despite being necessary, convergence to a weakly Pareto stationary point is thus not a sufficient condition for optimality and, hence, constitutes a rather weak guarantee. To the best of our knowledge, $\mathcal{A}_{\text{UPGrad}}$ is the first non-conflicting aggregator that provably converges to the Pareto front in the smooth convex case.
Appendix B.4 provides additional results and discussions on the convergence rate and about guarantees in the non-convex setting.
@@ 4 Existing aggregators
In the context of multi-task learning, several works have proposed iterative optimization algorithms based on the combination of task-specific gradients (Sener & Koltun, 2018; Yu et al., 2020; Liu et al., 2021b, a; Lin et al., 2021; Navon et al., 2022; Senushkin et al., 2023). These methods can be formulated as variants of JD parameterized by different aggregators. More specifically, since the gradients are stochastically estimated from batches of data, these are cases of what we call SEJD. In the following, we briefly present the most prominent aggregators and summarize their properties in Table 1. As a baseline, we also consider $\mathcal{A}_{\text{Mean}}$, which simply averages the rows of the Jacobian. Their formal definitions are provided in Appendix C. Some of them are also illustrated in Figure 1(b).
$\mathcal{A}_{\text{RGW}}$ aggregates the matrix using a random vector of weights (Lin et al., 2021). $\mathcal{A}_{\text{MGDA}}$ gives the aggregation that maximizes the smallest improvement (Fliege & Svaiter, 2000; Désidéri, 2012; Sener & Koltun, 2018). $\mathcal{A}_{\text{CAGrad}}$ maximizes the smallest improvement in a ball around the average gradient whose radius is parameterized by $c\in\left[0,1\right[{}$ (Liu et al., 2021a). $\mathcal{A}_{\text{PCGrad}}$ projects each gradient onto the orthogonal hyperplane of other gradients in case of conflict, iteratively and in a random order (Yu et al., 2020). It is, however, only non-conflicting when $m\varleq 2$, in which case $\mathcal{A}_{\text{PCGrad}}=m\cdot\mathcal{A}_{\text{UPGrad}}$. IMTL-G is a method to balance some gradients with impartiality (Liu et al., 2021b). It is only defined for linearly independent gradients, but we generalize it as a formal aggregator, denoted $\mathcal{A}_{\text{IMTL-G}}$, in Appendix C.6. Aligned-MTL orthonormalizes the Jacobian and weights its rows according to some preferences (Senushkin et al., 2023). We denote by $\mathcal{A}_{\text{Aligned-MTL}}$ this method with uniform preferences. $\mathcal{A}_{\text{Nash-MTL}}$ aggregates Jacobians by finding the Nash equilibrium between task-specific gradients (Navon et al., 2022). Lastly, the GradDrop layer (Chen et al., 2020) defines a custom backward pass that combines gradients with respect to some internal activation. The corresponding aggregator, denoted $\mathcal{A}_{\text{GradDrop}}$, randomly drops out some gradient coordinates based on their sign and sums the remaining ones.
@@
Several other works consider the gradients to be noisy when making their theoretical analysis (Zhou et al., 2022; Fernando et al., 2023; Chen et al., 2023; Xiao et al., 2023; Liu & Vicente, 2024). Their solutions for combining gradients are typically stateful. Although this could enhance practical convergence rates, we have restricted our focus to the analysis of stateless aggregators. Exploring and analyzing a generalized Jacobian descent algorithm, that would preserve some state over the iterations, is a promising future direction.
In the federated learning setting, several aggregators have been proposed to combine the model updates while being robust to adversaries (Blanchard et al., 2017; Chen et al., 2017; Guerraoui et al., 2018; Yin et al., 2018). We do not study them here as they mainly focus on security aspects.
Table 1: Properties satisfied for any number of objectives. Proofs are provided in Appendix C.
@@ 5 Experiments
Appendix A.1 shows the optimization trajectories of the methods from Table 1 when optimizing two convex quadratic forms that illustrate the setting of Figure 1 and the discrepancy between weak Pareto stationarity and Pareto optimality.
@@
We train convolutional neural networks on subsets of SVHN (Netzer et al., 2011), CIFAR-10 (Krizhevsky et al., 2009), EuroSAT (Helber et al., 2019), MNIST (LeCun et al., 1998), Fashion-MNIST (Xiao et al., 2017) and Kuzushiji-MNIST (Clanuwat et al., 2018). To make the comparisons as fair as possible, we have tuned the learning rate very precisely for each aggregator, as explained in detail in Appendix D.1. We have also run the same experiments several times independently to gain confidence in our results. Since this leads to a total of $43776$ training runs across all of our experiments, we have limited the size of each training dataset to 1024 images, greatly reducing computational costs. Note that this is strictly an optimization problem: we are not studying the generalization of the model, which would be captured by some performance metric on a test set. Other experimental settings, such as the network architectures and the total computational budget used to run our experiments, are given in Appendix D. Figure 2 reports the main results on SVHN and CIFAR-10, two of the datasets exhibiting the most substantial performance gap. Results on the other datasets and aggregators are reported in Appendix A.2. They also demonstrate a significant performance gap.
@@
The sub-optimal performance of $\mathcal{A}_{\text{MGDA}}$ in this setting can be attributed to its sensitivity to small gradients. If any row of the Jacobian approaches zero, the aggregation by $\mathcal{A}_{\text{MGDA}}$ will also approach zero. This observation illustrates the discrepancy between weak stationarity and optimality, as discussed in Section 2.4. A notable advantage of linearity under scaling is to explicitly prevent this from happening.
@@
While increasing the batch size in SGD reduces variance, the effect of doing so in SSJD combined with $\mathcal{A}_{\text{UPGrad}}$ is non-trivial, as it also tightens the dual cone. Additional results obtained when varying the batch size or updating the parameters with the Adam optimizer are available in Appendices A.3 and A.4, respectively.
While an iteration of SSJD is more expensive than an iteration of SGD, its runtime is influenced by several factors, including the choice of aggregator, the parallelization capabilities of the hardware used for Jacobian computation, and the implementation. Appendix E provides memory usage and computation time considerations for our methods. Additionally, we propose a path towards a more efficient implementation in the next section.
@@ 7 Conclusion
Our empirical experiments on deep learning problems have some limitations. First, we only evaluate JD on IWRM, a setting with moderately conflicting objectives. It would be essential to develop proper benchmarks to compare aggregators on a wide variety of problems. Ideally, such problems should involve substantially conflicting objectives, e.g. multi-task learning with inherently competing or even adversarial tasks. Then, we have limited our scope to the comparison of optimization speeds, disregarding generalization. While this simplifies the experiments and makes the comparison rigorous, optimization and generalization are sometimes intertwined. We thus believe that future works should focus on both aspects.
@@ Acknowledgments
Appendix A Additional experimental results
In this appendix, we provide additional experimental results about Jacobian descent and IWRM.
A.1 Optimization trajectories
Figures 3 and 4 show the optimization trajectories of the aggregators of Table 1 for two different convex quadratic functions, with several initializations [^4] We do not show the trajectories of $\mathcal{A}_{\text{PCGrad}}$ because it is equivalent to $2\cdot\mathcal{A}_{\text{UPGrad}}$ when there are only two objectives..
The first function is the element-wise quadratic function ${\bm{f}}_{\mathrm{EWQ}}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$, defined as
${\bm{f}}_{\mathrm{EWQ}}({\bm{x}})=\begin{bmatrix}x_{1}^{2}&x_{2}^{2}\end{bmatrix}^{\top}.$ (13)
Notably, its Pareto set only contains the origin, while its set of weakly Pareto stationary points is the union of the two axes. The Jacobian of ${\bm{f}}_{\mathrm{EWQ}}$ at $\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$ is $\begin{bmatrix}2x_{1}&0\\ 0&2x_{2}\end{bmatrix}$, indicating that the gradients never conflict, which makes this function relatively simple to optimize. Nevertheless, Figure 3 shows that $\mathcal{A}_{\text{MGDA}}$, $\mathcal{A}_{\text{CAGrad}}$, $\mathcal{A}_{\text{Nash-MTL}}$ and $\mathcal{A}_{\text{Aligned-MTL}}$ fail to converge to the Pareto set for some initializations. Indeed, they converge to weakly Pareto stationary points, which can be arbitrarily far from the Pareto set. This trivial example shows the importance of rather proving convergence to the Pareto front as in our Section 2.4.
The second function is a convex quadratic form ${\bm{f}}_{\mathrm{CQF}}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$, constructed to sometimes reproduce the setting of Figure 1 with conflicting and imbalanced gradients:
${\bm{f}}_{\mathrm{CQF}}\left({\bm{x}}\right)$ $=\begin{bmatrix}({\bm{x}}-{\bm{v}}_{1})^{\top}\cdot U\Sigma_{1}U^{\top}\cdot({\bm{x}}-{\bm{v}}_{1})\\ ({\bm{x}}-{\bm{v}}_{2})^{\top}\cdot U^{\top}\Sigma_{2}U\cdot({\bm{x}}-{\bm{v}}_{2})\end{bmatrix}$ (14)
with $\text{with }U$ $=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\text{, }\Sigma_{1}=\begin{bmatrix}1&0\\ 0&{\epsilon}\end{bmatrix}\text{, }\Sigma_{2}=\begin{bmatrix}3&0\\ 0&{\epsilon}\end{bmatrix}\text{, }\theta=\frac{\pi}{16}\text{, }{\epsilon}=0.01$ (15)
${\bm{v}}_{1}$ $=\begin{bmatrix}1\\ 0\end{bmatrix}\text{, and }{\bm{v}}_{2}=\begin{bmatrix}-1\\ 0\end{bmatrix}$ (16)
Figure 4 shows that most existing aggregators have some unusual behavior in this setting. In particular, on Figure 4(b), we can see that with $\mathcal{A}_{\text{Mean}}$, $\mathcal{A}_{\text{GradDrop}}$, $\mathcal{A}_{\text{IMTL-G}}$, $\mathcal{A}_{\text{CAGrad}}$, $\mathcal{A}_{\text{RGW}}$ and $\mathcal{A}_{\text{Aligned-MTL}}$, the updates can make one objective’s value increase. In contrast, with the non-conflicting aggregators $\mathcal{A}_{\text{MGDA}}$, $\mathcal{A}_{\text{DualProj}}$, $\mathcal{A}_{\text{Nash-MTL}}$ and $\mathcal{A}_{\text{UPGrad}}$, no objective value increases during the optimization. Due to its linearity under scaling, $\mathcal{A}_{\text{UPGrad}}$ also shows improved smoothness compared to the other non-conflicting aggregators.
(a) Parameter space. The black star is the Pareto set and the dashed lines are contour lines of the mean objective.
(b) Objective value space. The black star is the Pareto front. The background color represents the distance to the Pareto front.
Figure 3: Optimization trajectories of various aggregators when optimizing ${\bm{f}}_{\mathrm{EWQ}}:\begin{bmatrix}x_{1}&x_{2}\end{bmatrix}^{\top}\mapsto\begin{bmatrix}x_{1}^{2}&x_{2}^{2}\end{bmatrix}^{\top}$ with JD. Colored dots represent the initial points. The trajectories start in red and evolve towards yellow.
(a) Parameter space. The dark line is the Pareto set. The background color represents the cosine similarity between the two gradients, between $-1$ (fully conflicting) in pink and $1$ (fully aligned) in green. For instance, the similarity is $-0.925$ at the initial point $\begin{bmatrix}0&0\end{bmatrix}^{\top}$ and is $-1$ on (and only on) the Pareto set.
(b) Objective value space. The dark line is the Pareto front. The background color represents the distance to the Pareto front.
Figure 4: Optimization trajectories of various aggregators when optimizing the convex quadratic form ${\bm{f}}_{\mathrm{CQF}}$ of (14) with JD. Colored dots represent initial parameter values. The trajectories start in red and evolve towards yellow.
A.2 All datasets and all aggregators
Figures 5, 6, 7, 8, 9 and 10 show the full results of the experiments described in Section 5 on SVHN, CIFAR-10, EuroSAT, MNIST, Fashion-MNIST and Kuzushiji-MNIST, respectively. For readability, the results are displayed on three different plots for each dataset. We always show $\mathcal{A}_{\text{UPGrad}}$ and $\mathcal{A}_{\text{Mean}}$ for reference. The exact experimental settings are described in Appendix D.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 5: SVHN results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 6: CIFAR-10 results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 7: EuroSAT results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 8: MNIST results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 9: Fashion-MNIST results.
(a) Training loss
(b) Update similarity to the SGD update
(c) Training loss
(d) Update similarity to the SGD update
(e) Training loss
(f) Update similarity to the SGD update
Figure 10: Kuzushiji-MNIST results.
A.3 Varying the batch size
(a) $\text{BS}=4$: Training loss
(b) $\text{BS}=4$: Update similarity to the SGD update
(c) $\text{BS}=16$: Training loss
(d) $\text{BS}=16$: Update similarity to the SGD update
(e) $\text{BS}=64$: Training loss
(f) $\text{BS}=64$: Update similarity to the SGD update
A.4 Compatibility with Adam
Figure 12 gives the results on CIFAR-10 and SVHN when using Adam rather than the SGD optimizer. Concretely, this corresponds to the Adam algorithm in which the gradient is replaced by the aggregation of the Jacobian. The learning rate is still tuned as described in Appendix D.1, but the other hyperparameters of Adam are fixed to the default values of PyTorch, i.e. $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. Because optimization with Adam is faster, the number of epochs for SVHN and CIFAR-10 is reduced to 20 and 15, respectively. While the performance gap is smaller with this optimizer, it is still significant and suggests that our methods are beneficial with other optimizers than the simple SGD. Note that this analysis is fairly superficial. The thorough investigation of the interplay between aggregators and momentum-based optimizers is a compelling future research direction.
(a) SVHN: Training loss
(b) SVHN: Update similarity to the SGD update
(c) CIFAR-10: Training loss
(d) CIFAR-10: Update similarity to the SGD update
Figure 12: Results with the Adam optimizer.
Appendix B Proofs
B.1 Preliminary theoretical results